# Breaking Language Barriers or Reinforcing Bias? A Study of Gender and Racial Disparities in Multilingual Contrastive Vision–Language Models

### v1

@@ Appendix B Appendix: Prompt Labels
In order to systematically probe each stereotype axis, we constructed a set of prompt templates that pair a descriptive adjective or category with the phrase “a photo of …”. These templates were translated into each target language (see Section 7 for details on translation via GPT o3) and back-validated to ensure consistency. We break them down by axis below.
@@
The communion prompts aim to elicit cooperative and warmth-related stereotypes. Each descriptor reflects either a positive communion trait (e.g., trustworthy, benevolent) or its negative counterpart (e.g., untrustworthy, threatening). By comparing model responses across these twelve descriptors, we gauge the amplification of communal stereotypes.
@@
To probe dominance and power dynamics, we use twelve agency descriptors that capture both high-agency traits (e.g., powerful, confident) and their low-agency counterparts (e.g., powerless, meek). This set helps reveal stereotypes around competence and status.
@@
The crime prompts include both protected-class demographic labels (e.g., Black_Male, Middle_Eastern_Female) and explicit criminality terms (e.g., thief, suspicious person), enabling measurement of biased associations between demographics and “harmful” roles.

### v3

@@ Appendix F Prompt Labels
In order to systematically probe each stereotype axis, we constructed a set of prompt templates that pair a descriptive adjective or category with the phrase “a photo of …”. We break them down by axis below.
@@
The communion prompts aim to elicit cooperative and warmth‐related stereotypes. Each descriptor reflects either a positive communion trait (e.g., trustworthy, benevolent) or its negative counterpart (e.g., untrustworthy, threatening). By comparing model responses across these twelve descriptors, we gauge the amplification of communal stereotypes.
@@
To probe dominance and power dynamics, we use twelve agency descriptors that capture both high‐agency traits (e.g., powerful, confident) and their low‐agency counterparts (e.g., powerless, meek). This set helps reveal stereotypes around competence and status.
@@
The crime prompts include both protected‐class demographic labels (e.g., Black_Male, Middle_Eastern_Female) and explicit criminality terms (e.g., thief, suspicious person), enabling measurement of biased associations between demographics and “harmful” roles.